I was driving along at 20 m/s , trying to change a CD and not watching where I was going. When I looked up, I found myself 44 m from a railroad crossing. And wouldn't you know it, a train moving at 30 m/s was only 56 m from the crossing. In a split second, I realized that the train was going to beat me to the crossing and that I didn't have enough distance to stop. My only hope was to accelerate enough to cross the tracks before the train arrived. If my reaction time before starting to accelerate was 0.42 s , what minimum acceleration did my car need for me to be here today writing these words?
Answer #1

answer) the time taken by the train to reach crossing=56/30m/s=1.867s

the distance travelled in 0.42 s=0.42s*20m/s=8.4m

now distance you are from the crossing=44-8.4m=35.6m

you waited 0,42s , so total time you have to pass crossing before the train=1.867-0.42=1.447s

we know the formula

s=ut+1/2at2

35.6=20*1.447+1/2*a*1.447s

35.6=28.94+1.0469a

a=6.66/1.0469=6.36m/s2

so the answer is 6.37m/s2 or 6.36m/s2 or 6.4m/s2